# Fully automated workflow for the design of patient-specific orthopaedic implants: application to total knee arthroplasty

### v2

@@ Abstract
Osteoarthritis affects about 528 million people worldwide, causing pain, swelling and stiffness in the joints. Arthroplasty is commonly performed to treat joint osteoarthritis, reducing pain and improving mobility. While arthroplasty has known several technical improvements, a significant share of patients are still unsatisfied with their surgery. Personalised arthroplasty improves surgical outcomes however current solutions require delays, making it difficult to integrate in clinical routine. We propose a fully automated workflow to design patient-specific implants. The proposed approach is presented for total knee arthroplasty, the most widely performed arthroplasty in the world nowadays.
The proposed pipeline first uses artificial neural networks to segment the proximal and distal extremities of the femur and tibia bones. Then the full bones are reconstructed using augmented statistical shape models, combining shape and landmarks information. Finally, a total of 77 morphological parameters are computed to design patient-specific implants. The developed workflow has been trained using 91 CT scans of lower limb. Its performance has been evaluated on 41 CT scans manually segmented, in terms of accuracy and execution time.
The workflow accuracy was $0.4\pm 0.2mm$ for the segmentation, 1.$2\pm 0.4mm.$ for the full bones reconstruction, and $2.8\pm 2.2mm$ for the anatomical landmarks determination. The custom implants fitted the patients’ anatomy with $0.6\pm 0.2mm$ accuracy. The whole process from segmentation to implants’ design lasted about 5 minutes.
The proposed workflow allows for a fast and reliable personalisation of knee implants, directly from the patient CT image without requiring any manual intervention. It allows the establishment of a patient-specific pre-operative planning for TKA in a very short time making it easily available for all patients. Combined with efficient implant manufacturing techniques, this solution could help answer the growing number of arthroplasties while reducing complications and improving the patients’ satisfaction.
@@ 1 Introduction
Osteoarthritis (OA) affects about 528 million people worldwide, causing pain, swelling and stiffness in the joints [1]. Total joint arthroplasty is commonly performed to treat OA, when the joint is too severely affected and palliative approaches can no longer alleviate pain or improve mobility. Joint replacement surgery has evolved significantly since its conception. It gained acceptance with ever-improving implant survivorship and now aims to provide a ‘forgotten joint’ to most patients. To do so, personalising joint replacement is the key solution to restore native joint kinematics, function, and perception [2]. Continuous research efforts are conducted in that sense, mostly for hip and knee replacements. Indeed, they are by far the most performed arthroplasties nowadays, followed by shoulder, elbow, foot/ankle and hand/wrist replacements [3, 4]. Although they provide excellent clinical results, a major problem in total knee arthroplasty (TKA) compared to total hip arthroplasty, remains the high number of patients who are still not satisfied with the results of the surgery [5].
The number of TKA procedures has been rising steadily since the 1990s in Europe and the USA, and this increase is forecast to continue for the next decades [6, 7]. For instance, in France, the number of primary TKA reached 102 655 in 2019, (+62% increase compared to 2009), and is expected to increase to more than 150 000 TKA in 2050 [6]. Likewise, 480 958 primary TKAs were performed in the USA in 2019 (+45% compared to 2009) and recent literature predicts more than 1 800 000 TKAs in 2050 [7]. With a revision rate varying between 2% and 12% [8, 3, 9, 10], the number of revision TKA is also expected to grow until 2050 and so even more rapidly than primary TKA [11]. Two main factors can partially explain this growth: the ageing population and the expansion of TKA indication to younger and less severe symptoms patients [6]. While in 1997 patients younger than 65 years old represented 25% of TKA, they now represent 40% [12]. These patients have longer life expectancy, and so have higher risks to require a revision. Moreover, TKA for low grade OA represents higher costs than TKA for severe OA [13]. Therefore, TKA will inevitably represent a huge burden for the future health care system [11]. To answer this increasing demand and reduce the need for revision, surgeries have to be made faster while being more reliable, and implants durability have to be improved.
In the early years of TKA, the main challenge was to improve the implant stability. Consequently, the mechanical alignment (MA) technique was introduced to simplify and standardise the operations [14]. MA aims for a neutral mechanical axis to maximise the longevity of the TKA prosthesis by balancing load distribution between the medial and lateral compartments, and reducing polyethylene wear, excessive load on bone, and loosening [15]. Furthermore, the development of assistive technological tools, such as navigation, robotics and patient-specific instrumentation (PSI), have helped improve the accuracy of implantation and resection cuts to achieve accurate MA [14].
@@
These developments participated to improve surgical outcomes as well as the satisfaction rate among TKA patients. Between 85% and 95% of them were satisfied with their surgery over the last decade, when the satisfaction rate varied between 75% and 89% in the 2000s [14, 16]. Between 5% and 15% of TKA patients are then still unsatisfied with their surgery outcomes, mainly because of their increased expectations regarding symptoms, physical function, quality of life, coping strategies, and longevity of implant [17]. Several causes can be identified as source of dissatisfaction: overhang of the femoral component causing pain (almost 27% of all clinically important pain [18]); abnormal biomechanics, related to the cruciate ligaments removal as well as to the lack of consideration of the patient anatomy, limiting the mobility and stability of the knee [19]; or implants early mechanical fail, due for most to positioning, alignment or fixation defects [20].
@@
Consequently, TKA’s main focus is now shifting from enhancing implant survivorship toward improving patient function, kinematics and satisfaction. Currently, alternative alignment methods for TKA are gaining interest, to better respect the wide range of normal anatomy of the knee and restore individual anatomy with a personalised joint replacement. Different techniques have been described to obtain more anatomical and kinematic alignment than the conventional MA. For instance, recent studies considered pre-arthritic individual knee alignment and proposed different versions of kinematic alignment (KA) [21]. These different KA techniques all aim at resurfacing the femorotibial joint, to match the implant geometry with the bony anatomy in order to restore the native pre-arthritic limb alignment and adjust the ligament balance [15, 22].
@@
With TKA patients being more active today than ever, performing TKA that mimic the natural knee is essential to the patients’ long-term satisfaction and survival [23]. To do so, kinematic alignment techniques can be used along with personalised implants. Implant customisation aims to correct the patient’s knee deformities while staying as close as possible to his anatomy [24]. Custom implants offer three features that are rarely attainable when using off-the-shelf (OTS) implants: (1) Optimisation of the implant-bone fit, to avoid overhang or under-coverage; (2) Decoupling of the patellofemoral and tibiofemoral compartment, to optimise patellofemoral and tibiofemoral kinematics independently; and (3) Restoration of native condylar curvature, to improve ligament balancing, mid-flexion stability and kinematics [25, 26]. Although there is still no consensus on the advantages of custom implants compared to OTS implants in terms of satisfaction and pain scores [27, 28, 29], custom TKA demonstrated significant benefits regarding over- and under- sizing, local tendon impingement, improved kinematics, lower complication rate, and facilitation of restoration of constitutional coronal alignment [27]. However, replicating the patient anatomy is not sufficient to ensure the implant stability and patient satisfaction, as the cruciate ligaments are removed during the surgery and can no longer stabilise the joint. Combining custom TKA with ‘personalised alignment’ is expected to improve patient-reported outcome measures compared to OTS TKA [30] and represents one of the current biggest challenges to perform personalised TKA [27].
@@
Custom TKA implants have the potential to greatly improve knee kinematics and patient knee functions compared to OTS TKA implants. However, further investigation is needed to make the custom TKA implant readily accessible for patients [23]. Some commercialised solutions are already available to obtain custom implants, such as Origin® custom TKA (Symbios, Yverdon-les-Bains, Switzerland) or iTotalTMCR G2 (ConforMIS Inc., Bedford, MA, US). Both solutions provide a personalised implant along with patient-specific instrumentation (PSI) designed based on a computed tomography (CT) scan of the patient’s lower limb. The implants and PSI are delivered in a ready-to-use box, thus reducing the stock of implants and instrumentation needed in situ. However, the implant customisation is based on manual and semi-automated methods, making the design and manufacturing process last for up to 8 weeks [26, 31, 30]. Such processes need to be automated to allow a simple and fast establishment of the pre-operative planning for an easy integration in clinical routine and make personalised TKA available for patients needing surgery within less than two months.
@@
Therefore, to the best of our best knowledge, we propose the first fully automated workflow to design patient-specific TKA implants from CT images.
@@ 2.1 Global workflow
We developped a pre-operative planning software that can be connected to a PACS (Picture Archiving and Communication System) for direct access to the patients’ data. The different steps of the planning can be performed automatically within the software: segmentation of the hip, knee and ankle joints, determination of key anatomical landmarks and design of patient specific implants. The whole workflow is easily traceable as the result of each step is saved directly within the PACS under the patient ID and complies with DICOM format. An overview of the global workflow is proposed in Figure 1.
Figure 1: Global workflow overview
@@ 2.2 Dataset
Approval for this study was obtained from the Medical Ethics Committee of the University Hospital of Brest that coordinated the project under the Kneemod trial (N°29BRC18.0235), and is registered in Clinical Trial (NCT04179812). Participants gave informed written consent. 132 anonymised CT scans of lower limbs (right and/or left) from 75 patients were collected. These scans were acquired from 3 different machines (Toshiba® Aquilon One, Siemens® Somatom Edge Plus, Siemens® Somatom Definition Edge) at the Brest University Hospital during the years 2020-2021.
The volumes were manually segmented and then cropped to create three separate datasets for each joint: hip, knee and ankle. The data was split into training (91 scans from 50 patients) and test sets (41 scans from 25 patients).
@@ 2.3 CT image segmentation
After being predicted by our U-Net models, the segmentation masks were refined through a sequence of nonlinear operations related to their shape (morphological filtering). Additionally, a watershed algorithm [34] was applied to ensure bone separation in cases where delineation between the femur and tibia was ambiguous.
Finally, the 2D segmentation mask stacks were post-processed to construct 3D meshes of the bones, using the marching cubes and smoothing algorithms [35].
@@ 2.4.1 Augmented SSM building and fitting
Two SSMs have been built, one for each bone. All left femurs and tibias were initially flipped around their longitudinal axis to obtain only right bones datasets. Both right femur and right tibia SSMs have been built following the workflow previously described in [36]. First, a virtual reference shape is computed to establish unbiased correspondence between the training data and then, a principal component analysis is performed on the training data in correspondence. The femur and tibia SSMs have been built from training dataset shapes. The SSMs were augmented by integrating information on anatomical landmarks to the model (see next section for more details).
To reconstruct the whole femur and tibia bones from their proximal and distal extremities, we developed a custom algorithm to fit the SSM to partial data. Firstly, the SSM mean shape is scaled by adjusting the SSM first principal component to match the partial data bounding box’s lengths. Secondly, the partial data (or bone extremities) is rigidly registered to the scaled SSM mean shape. Thirdly, the SSM is deformed by optimising the coefficients along its principal components using a L-BFGS optimization [37]. Finally, a very coarse B-Spline surface is applied on the deformed SSM, to fit the bone surface as much as possible without reproducing osteophytes.
2.4.2 Extraction of knee morphometric data
To further design the custom implant, several anatomical landmarks were automatically identified.
The anatomical landmarks of the deformed SSM (inherited from the landmarks identified on the SSM mean shape) were used as initialisation for the landmarks detection. The landmarks were then adjusted to the segmented mesh to improve the detection accuracy. This adjustment was weighted regarding the risk of having osteophytes in the concerned zone.
A total of 77 morphological parameters were determined (48 femoral and 29 tibial). 23 primary anatomical landmarks were directly detected with the fitted SSMs (16 femoral and 7 tibial landmarks), and 54 secondary parameters (anatomical landmarks, axes, planes, pointset, lengths or angles) were computed from these primary landmarks. An example of secondary parameters computation is given in Figure 2 for the distal resection plane.
Figure 2: Computation flow of the femoral distal resection plane. Primary landmarks (left-most column) allow the computation of several secondary landmarks : anatomical points (purple boxes), axis (green boxes), sphere centres (blue) and planes (yellow boxes).
@@ 2.4.3 Validation
The quality of both tibia and femur models were evaluated by computing three metrics: compactness (ability to represent the variability of the training dataset), generality (ability to fit the testing dataset) and specificity (ability to generate shapes similar to the training dataset), as described by Davies et al. [38].
The full bone reconstruction accuracy has been tested by fitting the SSMs to the proximal and distal extremities of the 41 tibias and femurs of the testing dataset. The fitting duration has been registered and its accuracy has been evaluated by comparing the fitted models to the full bones manually segmented in terms of RMSE and HD.
The landmarks computation was validated based on the analysis of 11 femoral landmarks detailed in Figure 3. The computed landmarks were compared to ground truth landmarks acquired from manual picking on the 41 shapes of the testing dataset. The 11 femoral landmarks were determined twice by one surgeon on 3 modalities : CT images, virtual 3D model, and 3D printed model. The impact of the modality used on the landmarks determination accuracy, as well as the intra observer precision have been studied in a previous publication [39]. It was found that the existing variability on the acquired landmarks did not impact the axis and planes derived from them, and that no modality was more reliable than another. Therefore, the ground truth for each landmark was defined as the barycenter of the corresponding landmarks picked manually on each modality. The computed landmarks accuracy was evaluated in terms of distance (in millimetres) to the ground truth landmarks.
Figure 3: Anatomical landmarks acquired for the morphometric analysis validation.

### v3

# Fully automated workflow for designing patient-specific orthopaedic implants: application to total knee arthroplasty

@@ Abstract
Background Osteoarthritis affects about 528 million people worldwide, causing pain and stiffness in the joints. Arthroplasty is commonly performed to treat joint osteoarthritis, reducing pain and improving mobility. Nevertheless, a significant share of patients remain unsatisfied with their surgery. Personalised arthroplasty was introduced to improve surgical outcomes however current solutions require delays, making it difficult to integrate in clinical routine. We propose a fully automated workflow to design patient-specific implants for total knee arthroplasty.
Methods The proposed pipeline first uses artificial neural networks to segment the femur and tibia proximal and distal extremities. Then the full bones are reconstructed using augmented statistical shape models, combining shape and landmarks information. Finally, 77 morphological parameters are computed to design patient-specific implants. The developed workflow has been trained on 91 CT scans and evaluated on 41 CT scans, in terms of accuracy and execution time.
Results The workflow accuracy was $0.4\pm 0.2mm$ for segmentation, $1.0\pm 0.3mm$ for full bone reconstruction, and $2.2\pm 1.5mm$ for anatomical landmarks determination. The custom implants fitted the patients’ anatomy with $0.9\pm 0.5mm$ accuracy. The whole process from segmentation to implants’ design lasted about 15 minutes.
Conclusion The proposed workflow performs a fast and reliable personalisation of knee implants, directly from a CT image without requiring any manual intervention. It allows the establishment of a patient-specific pre-operative planning in a very short time, making it easily available for all patients. Combined with efficient implant manufacturing techniques, this solution could help answer the growing number of arthroplasties while reducing complications and improving patients’ satisfaction.
@@ 1 Introduction
Osteoarthritis (OA) affects about 528 million people worldwide, causing pain, swelling and stiffness in the joints [1]. Total joint arthroplasty is commonly performed to treat OA, when the joint is too severely affected and palliative approaches can no longer alleviate pain or improve mobility. Joint replacement surgery has evolved significantly since its conception. It gained acceptance with ever-improving implant survivorship and now aims to provide a ‘forgotten joint’ to most patients. To do so, personalising joint replacement is the key solution to restore native joint kinematics, function, and perception [2]. Continuous research efforts are conducted in that sense, mostly for hip and knee replacements. Indeed, they are by far the most performed arthroplasties nowadays, followed by shoulder, elbow, foot/ankle and hand/wrist replacements [3, 4]. Although total knee arthroplasty (TKA) can provide excellent clinical results, it suffers from a high number of patients who remain dissatisfied with the results of the surgery compared with total hip arthroplasty[5].
@@
On a first hand, the main challenge of TKA in its early years, was to improve the implant stability. Consequently, surgical techniques such as the mechanical alignment (MA) were introduced to simplify and standardise the operations [6]. MA also aimed to maximise the longevity of the TKA prosthesis by balancing load distribution between the implants’ compartments [7]. The development of assistive technological tools, such as navigation, robotics and patient-specific instrumentation, improved the accuracy of implantation with respect to the MA strategy, enhancing surgical outcomes as well as the satisfaction rate among TKA patients. Between 85% and 95% of them were satisfied with their surgery over the last decade, when the satisfaction rate varied between 75% and 89% in the 2000s [6, 8]. Then, between 5% and 15% of TKA patients are still unsatisfied with their surgery outcomes, mainly because of their increased expectations regarding symptoms, physical function, quality of life, coping strategies, and longevity of implant [9]. Several causes can be identified as source of dissatisfaction: overhang of the femoral component causing pain (almost 27% of all clinically important pain [10]); abnormal biomechanics, related to the cruciate ligaments removal as well as to the lack of consideration of the patient anatomy, limiting the mobility and stability of the knee [11]; or implants early mechanical fail, due for most to positioning, alignment or fixation defects [12]. In the end, such poor post-operative outcomes are likely to require revision surgery.
On a second hand, the number of TKA procedures has been rising steadily since the 1990s in Europe and the USA, and this increase is forecast to continue for the next decades [13, 14]. For instance, recent predictions for 2050 expect 150 000 TKAs in France (versus 102 655 in 2019), and more than 1 800 000 TKAs in the USA (versus 480 958 in 2019)[13, 14]. With a revision rate varying between 2% and 12% [15, 3, 16, 17], the number of revision TKA is also expected to grow until 2050 and so even more rapidly than primary TKA [18]. Two main factors can partially explain this growth: the ageing population and the expansion of TKA indication to younger and less severe symptoms patients [13]. While in 1997 patients younger than 65 years old represented 25% of TKA, they now represent 40% [19]. These patients have longer life expectancy, and so have higher risks to require a revision. Moreover, TKA for low grade OA represents higher costs than TKA for severe OA [20]. Therefore, TKA will inevitably represent a huge burden for the future health care systems [18].
@@
To answer TKA increasing demand and reduce the need for revision, surgeries have to be made faster while being more reliable. Consequently, TKA’s main focus is now shifting from enhancing implant survivorship toward improving patient function, kinematics and satisfaction. Currently, alternative techniques to MA are gaining interest, such as anatomical or kinematic alignment, to better respect the wide range of normal anatomy of the knee and restore individual anatomy with a personalised joint replacement [21]. These alignment techniques all aim at resurfacing the femorotibial joint, matching the implant geometry with the bony anatomy to restore the native pre-arthritic limb alignment and adjust the ligament balance [7, 22].
@@
With TKA patients being more active today than ever, performing TKA that mimic the natural knee is therefore essential to the patients’ long-term satisfaction and survival [23]. To restore kinematic function, implant customisation aims to correct the patient’s knee deformities while staying as close as possible to his anatomy [24, 25]. Custom implants offer three features that are rarely attainable when using off-the-shelf (OTS) implants: (1) Optimisation of the implant-bone fit, to avoid overhang or under-coverage; (2) Decoupling of the patellofemoral and tibiofemoral compartment, to optimise patellofemoral and tibiofemoral kinematics independently; and (3) Restoration of native condylar curvature, to improve ligament balancing, mid-flexion stability and kinematics [26, 27]. Although there is still no consensus on the advantages of custom implants compared to OTS implants in terms of satisfaction and pain scores [28, 29, 30], custom TKA demonstrated significant benefits regarding over- and under- sizing, local tendon impingement, improved kinematics, lower complication rate, and facilitation of restoration of constitutional coronal alignment [28]. However, replicating the patient anatomy is not sufficient to ensure the implant stability and patient satisfaction, as the cruciate ligaments are removed during the surgery and can no longer stabilise the joint. Combining custom TKA with latest prosthetic concepts and ‘personalised alignment’ is expected to improve patient-reported outcome measures compared to OTS TKA [31] and represents one of the current biggest challenges to perform personalised TKA [28].
@@
Custom TKA implants have the potential to greatly improve knee kinematics and patient knee functions compared to OTS TKA implants. However, further investigation is needed to make the custom TKA implant readily accessible for patients [23]. Today, only two solutions are proposed on the market of personalised implants: Origin® custom TKA (Symbios, Yverdon-les-Bains, Switzerland) and iTotalTMCR G2 (ConforMIS Inc., Bedford, MA, US). Both solutions provide a personalised implant along with patient-specific instrumentation designed based on a computed tomography (CT) scan of the patient’s lower limb. The personalised implants and instrumentation are delivered in a ready-to-use box, thus reducing the stock of implants and instrumentation needed in situ. However, the implant customisation is based on manual and semi-automated methods, making the design and manufacturing process last for up to 8 weeks, from planning to delivering the final implant [27, 25, 31]. Furthermore, establishing the pre-operative planning and designing custom implants require several back and forth between surgeons and the engineering team, making it difficult to personalise surgery on a day-to-day basis. Such processes need to be automated to allow a simple and fast establishment of the pre-operative planning for a seamless integration in clinical routine[32] and make personalised TKA available for patients needing surgery within less than two months. Recent research efforts have been undertaken to automatise implants personalisation, but are limited by one or more of the following factors: they have focused solely on the femoral component, did not consider knee alignment, have been tested on a single study case, or have involved semi automated or manual processes[33, 34].
@@
Therefore, to the best of our knowledge, we propose the first fully automated workflow to design patient-specific TKA implants from CT images.
@@ 2.1 Global workflow
We developed a pre-operative planning software that can be connected to a PACS (Picture Archiving and Communication System) for direct access to the patients’ data. The different steps of the planning can be performed automatically within the software: segmentation of the hip, knee and ankle joints, determination of key anatomical landmarks and design of patient-specific implants. The whole workflow is easily traceable, as the result of each step is saved directly within the PACS under the patient ID and complies with DICOM format. An overview of the global workflow is proposed in Fig 1.
Fig 1: Global workflow overview to generate patient-specific orthopaedic implants.
@@ 2.2 Dataset
132 anonymised CT scans of lower limbs (right and/or left) from 75 patients were collected on 22/10/2021. These scans were acquired from 3 different machines (Toshiba® Aquilon One, Siemens® Somatom Edge Plus, Siemens® Somatom Definition Edge) at the Brest University Hospital during the years 2020-2021. Approval for this study was obtained from the Medical Ethics Committee of the University Hospital of Brest that coordinated the project under the Kneemod trial (N°29BRC18.0235), and is registered in Clinical Trial (NCT04179812). Participants gave informed written consent. Patients were included if CT images of their leg were present in the University Hospital of Brest database, they had knee trauma, osteoarthritic or healthy knee, and they formulated non opposition. Patients were excluded if they refused to participate or if only partial views were available, thus not allowing a good segmentation of the total joints. The volumes were manually segmented and then cropped to create three separate datasets for each joint: hip, knee and ankle. The data was split into training (91 scans from 50 patients) and test sets (41 scans from 25 patients).
@@ 2.3 CT image segmentation
The predicted 2D segmentation masks were post-processed with morphological filtering and a watershed algorithm [37] to ensure femur and tibia separation in ambiguous cases. Finally, the 2D segmentation masks were stacked to construct 3D meshes of the bones, using the marching cubes and smoothing algorithms [38].
@@ 2.4.1 Augmented SSM building and fitting
Two SSMs have been built, one for each bone, from the training dataset. All left femurs and tibias were initially flipped around their longitudinal axis to obtain only right bones datasets. Both femur and tibia SSMs have been built following the workflow previously described in [39]. First, a virtual reference shape is computed to establish unbiased correspondence between the training data. Then, a principal component analysis is performed on the training data in correspondence to compute the SSM’s mean shape and modes of deformation. The SSMs were augmented by integrating information on anatomical landmarks to the model (see next section for more details).
To reconstruct the whole femur and tibia bones from their proximal and distal extremities, a custom algorithm was developed to fit a SSM to partial data. Firstly, the SSM mean shape is scaled to match the partial data bounding box’s lengths. Secondly, the scaled SSM mean shape is rigidly registered to the partial data. Thirdly, the SSM is deformed to fit the bone surface as much as possible without reproducing osteophytes.
2.4.2 Extraction of morphometric data
To further design the custom implant, several anatomical landmarks were automatically identified. The anatomical landmarks of the deformed SSM (inherited from the landmarks identified on the SSM mean shape) were used as initialisation for the landmarks detection. The landmarks were then adjusted to the segmented mesh to improve the detection accuracy. This adjustment was weighted regarding the risk of having osteophytes in the concerned zone.
A total of 77 morphological parameters (48 femoral and 29 tibial) were determined on the bone 3D models. 23 primary anatomical landmarks were directly detected with the fitted SSMs (16 femoral and 7 tibial landmarks), and 54 secondary parameters (anatomical landmarks, axes, planes, pointset, lengths or angles) were computed from these primary landmarks.
@@ 2.4.3 Validation
The quality of both tibia and femur SSMs were evaluated by computing three metrics, as described by Davies et al. [40]: compactness (ability to represent the variability of the training dataset), generality (ability to fit the testing dataset) and specificity (ability to generate shapes similar to the training dataset).
The full bone reconstruction accuracy was assessed by fitting the SSMs to the proximal and distal extremities of the 41 tibias and femurs of the testing dataset. The fitting duration has been registered and its accuracy has been evaluated by comparing the fitted models to the full bones manually segmented in terms of RMSE and HD.
The landmarks computation was validated based on the analysis of 17 primary landmarks (11 femoral and 6 tibial) detailed in Fig 2. The computed landmarks were compared to ground truth landmarks acquired manually on the 41 shapes of the testing dataset. The 11 femoral landmarks were determined twice by two orthopaedic surgeons on 3 modalities : CT images, virtual 3D models, and 3D printed models. A previous study showed that no modality was more reliable than another and that the intra observer variability did not impact the computation of secondary parameters [41]. The 7 tibial landmarks were thus determined twice by two orthopaedic surgeons on virtual 3D models only. The ground truth for each landmark was defined as the barycenter of the corresponding landmarks picked manually. The computed landmarks accuracy was evaluated in terms of distance (in millimetres) to the ground truth landmarks. As an exception, the tibial condyles centres were computed from 5 points picked on the edge of both the lateral and medial plateau, and their accuracy was evaluated through the angular error (in degrees) of the line they form.
Fig 2: Femoral and tibial landmarks acquired for the morphometric analysis validation. The condyles centers were determined from 5 points acquired on the edge of the lateral and medial plateau (blue lines).